Circular No. 6455

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1996 Q1 Vello Tabur, Wanniassa, Australian Capital Territory, reports his dis-

covery of a comet. The following observations are available:

1996 UT  α_{2000}  δ_{2000}  m_{1} Observer
Aug. 19.6951 4^{h}44^{m}30^{s}  −9^{o}23.^{′}2  10  Tabur
19.8000 4 44 42  −9 22.4  _{”}
20.7361 4 46 12  −9 02.9  _{”}
21.10556 4 46 44.91 −8 55 52′′.2 11.1 Mikuž
21.11506 4 46 45.85 −8 55 41.1  _{”}
21.12133 4 46 46.38 −8 55 36.9  _{”}
21.12574 4 46 46.78 −8 55 28.3  _{”}
21.45 4 47.5 −8 49 10.0 Hale

V. Tabur (Wanniassa, A.C.T.). 0.20-m f/4.7 reflector. Circular coma of

diameter 3′ with little or no condensation. H. Mikuž (Crni Vrh Observatory). 0.20-m f/2 Baker-Schmidt telescope +

CCD + V filter. Coma diameter 4′.5, moderately condensed. A. Hale (Cloudcroft, NM). 0.41-m reflector (visual). Coma diameter 3′.5.

GRS 1915+105 B. Sams and A. Eckart, Max-Planck-Institut für Extraterrestrische

Physik; and R. Sunyaev, Max-Planck-Institut für Astrophysik, report the disappearance of the near-infrared jets observed in GRS 1915+105 in July 1995 (Sams et al. 1996, Nature 382, 47): “Using the same observing equipment as in July 1995 (ESO New Technology Telescope + SHARP speckle camera), new observations on June 7 show no evidence of any asymmetry in the image of the object down to a 3σ level of (magnitude) K = 16.5 ± 0.3. This is a factor of eleven times fainter than the previously observed southwest jet. Hence the lifetime of the infrared jet is < 1 yr. We detected the brightest total infrared emission yet observed from GRS 1915+105, with J = 16.3 ± 0.1, H = 14.0 ± 0.1, and K = 12.3 ± 0.1. Fluxes were calibrated against κ Aql (HR 7446), which was assumed to have J = 5.03, H = 5.06, and K = 5.07. Comparison with previous observations (IAUC 6267; Mirabel et al. 1995, A.Ap. 282, L17) shows that the infrared emission from GRS 1915+105 becomes bluer as the K-band emission increases. This behavior is typical of systems with x-ray heating of the normal star and outer zones of the accretion disk.”

1996 August 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green